OpenStudy (anonymous):

What is the solution to 2x^2+x+2=0?

OpenStudy (solomonzelman):

what methods do you know? factoring completing the square quadratic formula

OpenStudy (charlotte123):

all I got was this >.> http://www.wolframalpha.com/input/?i=2x%5E2%2Bx%2B2%3D0

OpenStudy (anonymous):

@SolomonZelman?

OpenStudy (solomonzelman):

@charlotte123, why are you posting a solving calc for this, this is just an alg 2 staff.

OpenStudy (anonymous):

this is the question

OpenStudy (anonymous):

7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n) these are the choices: 1 3/5, 3/8, 5/14, 6 6/7

OpenStudy (solomonzelman):

Wow, not here.

OpenStudy (anonymous):

ok where

OpenStudy (solomonzelman):

@mrivera3113, can you solve it?

OpenStudy (anonymous):

-1 -+ I sqroot of 15 / 4

OpenStudy (solomonzelman):

 this is what I got.